# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ Abstract
This paper aims to conduct a comprehensive theoretical analysis of current diffusion models. We introduce a novel generative learning methodology utilizing the Schrödinger bridge diffusion model in latent space as the framework for theoretical exploration in this domain. Our approach commences with the pre-training of an encoder-decoder architecture using data originating from a distribution that may diverge from the target distribution, thus facilitating the accommodation of a large sample size through the utilization of pre-existing large-scale models. Subsequently, we develop a diffusion model within the latent space utilizing the Schrödinger bridge framework. Our theoretical analysis encompasses the establishment of end-to-end error analysis for learning distributions via the latent Schrödinger bridge diffusion model. Specifically, we control the second-order Wasserstein distance between the generated distribution and the target distribution. Furthermore, our obtained convergence rates effectively mitigate the curse of dimensionality, offering robust theoretical support for prevailing diffusion models.
@@ 1 Introduction
In this paper, we advance the theoretical analysis of diffusion models with latent space by incorporating a thorough exploration of the encoder-decoder structure. We propose a novel latent diffusion model based on the Schrödinger bridge [Sch32, Jam75, DP91, Léo14] as the framework for theoretical exploration in diffusion models. Currently, a distinct line of inquiry in diffusion models has emerged, rooted in the Schrödinger bridge [Sch32, Jam75, DP91, Léo14], substantiated by pertinent literature references [WJX+21, DBTHD21, HHLP23, LVH+23, CHZ+23], wherein the SDEs are defined in a finite time horizon. This differs from diffusion models predominantly based on [SSDK+20], wherein the adopted SDEs, such as Ornstein-Uhlenbeck (OU) and Langevin SDEs, are defined over an infinite time horizon $[0,\infty)$. The diffusion models grounded in the Schrödinger bridge, as evidenced by [WJX+21, DBTHD21, HHLP23, LVH+23, CHZ+23], have demonstrated exceptional performance in practical applications, surpassing existing diffusion models. Additionally, from both sampling and optimization perspectives [HJK+21, JKLZ21, DJK+23], the Schrödinger bridge has demonstrated its superiority over Langevin SDE methods. Furthermore, in [WJX+21], a convergence analysis is conducted for the Schrödinger bridge generative model, and [DBTHD21] provided a theoretical analysis under an accurate score estimation and some regularity assumptions. However, these assumptions warrant further verification, and their frameworks do not consider the pre-training encoder-decoder structure. In our method, we first train an encoder-decoder structure separately, a process which can be conducted with domain shift. Then, we construct an SDE defined over the time interval $[0,1]$ in the latent space, where the initial distribution is the convolution of the encoder target distribution and a Gaussian distribution, and the terminal distribution is the encoder target distribution. Thirdly, we derive the deep score estimation of the score function corresponding to the initial convolution distribution using score matching techniques [HD05, Vin11]. Finally, we employ the Euler–Maruyama (EM) approach to discretize the SDE corresponding to the estimated score, thereby obtaining the desired samples by implementing the early stopping technique and the trained decoder. See Section 3 for a detailed description of our proposed latent diffusion models. It is imperative to acknowledge that within our method, a salient observation arises regarding the decoupling of the pre-training procedure from the formulation of the latent diffusion model. Moreover, the pre-training process is endowed with the capacity to accommodate distributional shifts diverging from the target distribution. Consequently, this approach facilitates the seamless incorporation of pre-existing large-scale models into the pre-training phase, thus guaranteeing an abundant supply of samples for training endeavors. Theoretically, we systematically establish rigorous end-to-end Wasserstein-distance error bounds between the distribution of generated data and the target distribution.
@@
  • Building upon the Schrödinger bridge, we introduce a novel latent diffusion model for generative learning. Our method incorporates a separate encoder-decoder structure, which can be utilized to accommodate domain shift. This can be accomplished by integrating pre-existing large-scale models into the pre-training phase, ensuring the availability of an ample corpus of samples essential for comprehensive pre-training procedures. We formulate an SDE defined on the time interval $[0,1]$, transitioning from the Gaussian-convoluted distribution of the encoder target distribution to the encoder target distribution. Leveraging score matching techniques [HD05, Vin11], EM discretization, early stopping technique, and the trained decoder, we derive samples that are approximately distributed according to the target distribution.
  • We rigorously establish the end-to-end theoretical analysis for latent diffusion models. Specifically, through a combination of pre-training, score estimation, numerical analysis of SDEs, and time truncation techniques, we deduce the Wasserstein-distance error bounds between the distribution of generated samples and the target distribution, quantified as $\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}\right)$. Here, $n$ denotes the sample size used to train latent diffusion model, and $d^{*}$ represents the dimension of latent space. This rate mitigates the curse of dimensionality inherent in raw data, thereby illustrating the effectiveness of latent diffusion models from a theoretical perspective. Moreover, the framework of our theoretical analysis is also applicable for analyzing other diffusion models.
@@ 1.1 Related work
Normalizing flows [DSDB16, RM15, PNR+21], especially those guided by ordinary differential equations (ODEs), provide a robust framework for transforming a simple probability distribution, such as a standard normal distribution, into a more complex form through a sequence of invertible and differentiable mappings. The essence of normalizing flows lies in their ability to capture complex dependencies and structures within data by iteratively applying invertible transformations. These transformations facilitate the modeling of sophisticated probability distributions, enabling practitioners to navigate through the intricacies of high-dimensional data and capture latent features. Recently, [CDD23, CCL+23] undertook a theoretical analysis of ODE flows, specifically introducing the accurate score estimation. However, this assumption warrants further verification. Very recently, [CDJ+24, GHJZ24, JLWY24] provide end-to-end convergence rates for ODE-based (conditional) generative models.
@@
In diffusion models, the fundamental architecture comprises both forward and backward processes. The primary objective of the forward process is to systematically add Gaussian noises to the dataset, thereby resulting in the limiting distribution being a Gaussian distribution. Conversely, the backward process assumes the role of a denoising mechanism, undertaking the formidable task of reconstructing data in a manner that closely approximates the characteristics of the original samples. Within the denoising step, a critical step emerges in the form of training a score neural network. This neural network is meticulously crafted to approximate the underlying score function, the gradient of the log density function of the underlying distribution. This synthesis of forward and backward processes, coupled with the sophisticated training of a score neural network, encapsulates the essence of diffusion models. Furthermore, within practical applications, datasets frequently exhibit inherent low-dimensional structures. The derivation of their low-dimensional representations within latent spaces is consequently pursued. A noteworthy advancement within this domain is the stable diffusion model, introduced by [RBL+22]. The integration of an autoencoder architecture within the stable diffusion model emerges as a pivotal mechanism, facilitating the discernment of the intrinsic low-dimensional essence of raw data. By executing diffusion processes within the latent spaces, the model excels in encapsulating the underlying distribution. This capability is particularly advantageous within complex data distributions. In addition, analogous studies have investigated the effective capture of data distributions through the utilization of latent spaces. Reference can be made to [VKK21, RDN+22]. In summary, these diffusion models can be generally classified into three categories: denoising diffusion probabilistic models, score-based generative models, and score SDEs. A thorough examination of diffusion models can be found in the specialized review paper [YZS+23]. The theoretical analysis of diffusion models primarily focuses on error analysis with accurate score estimators [CLL23, CDS23, LLT22, LLT23, BDBDD23, LWCC23, GNZ23] or end-to-end error analysis [OAS23, CHZW23]. However, these theoretical results did not consider the theoretical analysis of pre-training encoder-decoder structures, a crucial step in contemporary mainstream diffusion models. Consequently, these findings are unable to provide further insights into the underlying mechanisms of diffusion models. In contrast, our theoretical results include the theoretical analysis of pre-training encoder-decoder architectures, thereby offering a comprehensive explanation.
@@ 1.2 Notations
We introduce the notations used throughout this paper. Let $[N]:=\{0,1,\cdots,N-1\}$ represent the set of integers ranging from 0 to $N-1$. Let $\mathbb{N}^{+}$ denote the set of positive integers. For matrices $A,B\in\mathbb{R}^{d\times d}$, we assert $A\preccurlyeq B$ when the matrix $B-A$ is positive semi-definite. The identity matrix in $\mathbb{R}^{d\times d}$ is denoted as $\mathbf{I}_{d}$. The $\ell^{2}$-norm of a vector $\mathbf{x}=\{x_{1},\ldots,x_{d}\}^{\top}\in\mathbb{R}^{d}$ is defined by $\|\mathbf{x}\|:=\sqrt{\sum_{i=1}^{d}x_{i}^{2}}$. Simultaneously, the operator norm of a matrix $A$ is articulated as $\|{A}\|:=\sup_{\|\mathbf{x}\|\leq{1}}\|{A\mathbf{x}}\|$. The function space $C^{2}(\mathbb{R}^{d})$ encompasses functions that are twice continuously differentiable from $\mathbb{R}^{d}$ to $\mathbb{R}$. For any $f\in C^{2}(\mathbb{R}^{d})$, the symbols $\nabla{f}$, $\nabla^{2}{f}$, and ${\rm{\Delta}}f$ signify its gradient, Hessian matrix, and Laplacian, respectively. The $L^{\infty}(K)$-norm, denoted as $\|{f}\|_{L^{\infty}(K)}:=\sup_{\mathbf{x}\in K}|f(\mathbf{x})|$, captures the supremum of the absolute values of a function over a set $K\subset\mathbb{R}^{d}$. For a vector function $\mathbf{v}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, the $L^{\infty}(K)$-norm is defined as $\|\mathbf{v}\|_{L^{\infty}(K)}:=\sup_{\mathbf{x}\in K}\|\mathbf{v}(\mathbf{x})\|$. The asymptotic notation $f(\mathbf{x})=\mathcal{O}\left(g(\mathbf{x})\right)$ is employed to signify that $f(\mathbf{x})\leq Cg(\mathbf{x})$ for some constant $C>0$. Additionally, the notation $\widetilde{\mathcal{O}}(\cdot)$ is utilized to discount logarithmic factors in the asymptotic analysis.
@@ 2 Preliminaries
A class of neural networks NN$(L,M,J,K,\kappa)$ with depth $L$, width $M$, sparsity level $J$, boundness $K$, weight $\kappa$, is defined as
@@
$\mathbf{W}_{i}\in\mathbb{R}^{d_{i+1}\times d_{i}},b_{i}\in\mathbb{R}^{d_{i+1}},i=0,1,\ldots,L-1,$
the width is $\text{the width is $M:=\max\{d_{0},\ldots,d_{L}\}$},$
$\mathop{\sup}_{t,\mathbf{x}}\|{\mathbf{s}(t,\mathbf{x})}\|\leq{K},\mathop{\max}_{1\leq{i}\leq{L}}\{\|{\mathbf{b}}_{i}\|_{\infty},\|{\mathbf{W}_{i}}\|_{\infty}\}\leq{\kappa},$
$\sum_{i=1}^{L}\left(\|{\mathbf{W}_{i}}\|_{0}+\|{\mathbf{b}_{i}}\|_{0}\right)\leq J\Big{\}}.$
@@ 3.1 Pre-training
We observe that the derivation of $(\widehat{\boldsymbol{E}},\widehat{\boldsymbol{D}})$ as defined in (3) can be facilitated through the utilization of pre-existing large-scale models. This approach permits accommodation for distributional shifts that may deviate from the target distribution. As a result, it guarantees an abundant supply of samples essential for pre-training endeavors, enabling us to obtain convergence rates that circumvent the curse of dimensionality, as discussed in our theoretical development in Section 4.2.4 and Section 4.3.
@@ 3.2 Latent Diffusion Modeling
With the trained encoder-decoder structure established in Section 3.1, we can subsequently construct the diffusion model in the latent space. We denote by $\widehat{\mathbf{x}}:=\widehat{\boldsymbol{E}}\mathbf{x}\in\mathbb{R}^{d^{*}}$ with $\mathbf{x}\sim p_{data}(\cdot)$. Let ${p}^{*}_{data}(\cdot)$ and $\widehat{p}^{*}_{data}(\cdot)$ represent the density functions of $\boldsymbol{E}^{*}\mathbf{x}$ and $\widehat{\mathbf{x}}$, respectively. Additionally, let $\Phi_{\sigma}(\cdot)$ denote the density of the normal distribution $\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{d^{*}})$. Let $\mathcal{X}:=\{\mathbf{x}_{i}\}_{i=1}^{n}$ be i.i.d. from $p_{data}$, then we have access to the encoder data $\widehat{\mathcal{X}}:=\{\widehat{\mathbf{x}}_{i}\}_{i=1}^{n}=\{\widehat{\boldsymbol{E}}\mathbf{x}_{i}\}_{i=1}^{n}$. We denote the convolution of $\widehat{p}^{*}_{data}$ and $\Phi_{\sigma}(\cdot)$ as $q(\cdot,\cdot)$, defined as follows:
@@
$\mathrm{d}\mathbf{x}_{t}=\sigma^{2}\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\mathrm{d}t+\sigma\mathrm{d}\mathbf{w}_{t},\leavevmode\nobreak\ t\in[0,1],\leavevmode\nobreak\ \mathbf{x}_{0}\sim{q}({\sigma},\mathbf{x}),$ (5)
where $\mathbf{w}_{t}$ represents a $d^{*}$-dimensional Brownian motion. Consequently, $\mathbf{x}_{1}$ follows the distribution $\widehat{p}^{*}_{data}$ as established in [WJX+21, Theorem 3]. This formalization provides a coherent and rigorous foundation for our generative framework.
Score Matching. The SDE provided in (5) offers a pathway for constructing a generative model through the application of EM discretization, contingent upon access to the target score denoted as $\mathbf{s}^{*}(t,\mathbf{x}):=\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})$. However, in practical scenarios, this score function remains elusive due to its intricate connection with the target density $\widehat{p}^{*}(\cdot)$. Consequently, we resort to score matching techniques [HD05, Vin11] employing deep ReLU FNNs.
Verification reveals that $s^{*}$ minimizes the loss function $\mathcal{L}(s)$ across all measurable functions, with $\mathcal{L}(s)$ expressed as
$\mathcal{L}(\mathbf{s}):=\frac{1}{T}\int_{0}^{T}\left(\mathbb{E}_{\mathbf{x}_{t}\sim q(\sqrt{1-t}\sigma,\mathbf{x})}\|{\mathbf{s}(t,\mathbf{x}_{t})-\nabla\log{q(\sqrt{1-t}\sigma,\mathbf{x}_{t})}}\|^{2}\right)\mathrm{d}t,$
where $10<T<$ 1. We notice that the deliberate selection of $T$ is motivated by the necessity to preclude the score function from exhibiting a blow-up at $T=1$, concomitantly facilitating the stabilization of the training process for the model. A meticulous exposition regarding the determination of $T$ is provided in Section 4.2.4. Furthermore, the application of this time truncation technique extends beyond our work and has found utility in both model training and theoretical analyses in diffusion models [SE20, VKK21, CHZW23, CLL23, OAS23].

### v2

@@ Abstract
This paper aims to conduct a comprehensive theoretical analysis of current diffusion models. We introduce a novel generative learning methodology utilizing the Schrödinger bridge diffusion model in latent space as the framework for theoretical exploration in this domain. Our approach commences with the pre-training of an encoder-decoder architecture using data originating from a distribution that may diverge from the target distribution, thus facilitating the accommodation of a large sample size through the utilization of pre-existing large-scale models. Subsequently, we develop a diffusion model within the latent space utilizing the Schrödinger bridge framework. Our theoretical analysis encompasses the establishment of end-to-end error analysis for learning distributions via the latent Schrödinger bridge diffusion model. Specifically, we control the second-order Wasserstein distance between the generated distribution and the target distribution. Furthermore, our obtained convergence rates are sharp and effectively mitigate the curse of dimensionality, offering robust theoretical support for prevailing diffusion models.
@@ 1 Introduction
In this paper, we advance the theoretical analysis of diffusion models with latent space by incorporating a thorough exploration of the encoder-decoder structure. We propose a novel latent diffusion model based on the Schrödinger bridge [Sch32, Jam75, DP91, Léo14] as the framework for theoretical exploration in diffusion models. Currently, a distinct line of inquiry in diffusion models has emerged, rooted in the Schrödinger bridge [Sch32, Jam75, DP91, Léo14], substantiated by pertinent literature references [WJX+21, DBTHD21, HHLP23, LVH+23, CHZ+23], wherein the SDEs are defined in a finite time horizon. This differs from diffusion models predominantly based on [SSDK+20], wherein the adopted SDEs, such as Ornstein-Uhlenbeck (OU) and Langevin SDEs, are defined over an infinite time horizon $[0,\infty)$. The diffusion models grounded in the Schrödinger bridge, as evidenced by [WJX+21, DBTHD21, HHLP23, LVH+23, CHZ+23], have demonstrated exceptional performance in practical applications, surpassing existing diffusion models. Additionally, from both sampling and optimization perspectives [HJK+21, JKLZ21, DJK+23], the Schrödinger bridge has demonstrated its superiority over Langevin SDE methods. Furthermore, in [WJX+21], a convergence analysis is conducted for the Schrödinger bridge generative model, and [DBTHD21] provided a theoretical analysis under an accurate score estimation and some regularity assumptions. However, these assumptions warrant further verification, and their frameworks do not consider the pre-training encoder-decoder structure. In our method, we first train an encoder-decoder structure separately, a process which can be conducted with domain shift. Then, we construct an SDE defined over the time interval $[0,1]$ in the latent space, where the initial distribution is the convolution of the encoder target distribution and a Gaussian distribution, and the terminal distribution is the encoder target distribution. Thirdly, we derive the deep score estimation of the score function corresponding to the initial convolution distribution using score matching techniques [HD05, Vin11]. Finally, we employ the Euler–Maruyama (EM) approach to discretize the SDE corresponding to the estimated score, thereby obtaining the desired samples by implementing the early stopping technique and the trained decoder. See Section 3 for a detailed description of our proposed latent diffusion models. It is imperative to acknowledge that within our method, a salient observation arises regarding the decoupling of the pre-training procedure from the formulation of the latent diffusion model. Moreover, the pre-training process is endowed with the capacity to accommodate distributional shifts diverging from the target distribution. Consequently, this approach facilitates the seamless incorporation of pre-existing large-scale models into the pre-training phase, thus guaranteeing an abundant supply of samples for training endeavors. Theoretically, we systematically establish rigorous end-to-end Wasserstein-distance error bounds between the distribution of generated data and the target distribution, which are minimax optimal.
@@
  • Building upon the Schrödinger bridge, we introduce a novel latent diffusion model for generative learning. Our method incorporates a separate encoder-decoder structure, which can be utilized to accommodate domain shift. This can be accomplished by integrating pre-existing large-scale models into the pre-training phase, ensuring the availability of ample samples essential for comprehensive pre-training procedures. We formulate an SDE defined on the time interval $[0,1]$, transitioning from the Gaussian-convoluted distribution of the encoder target distribution to the encoder target distribution. Leveraging score matching techniques [HD05, Vin11], EM discretization, early stopping technique, and the trained decoder, we derive samples that are approximately distributed according to the target distribution.
  • We rigorously establish the end-to-end theoretical analysis for latent diffusion models. Specifically, through a combination of pre-training, score estimation, numerical analysis of SDEs, and time truncation techniques, we deduce the optimal Wasserstein-distance error bounds between the distribution of generated samples and the target distribution, quantified as $\widetilde{\mathcal{O}}\left(n^{-\frac{\beta}{d^{*}+2\beta}}\right)$. Here, $n$ denotes the sample size used to train latent diffusion model, $d^{*}$ represents the dimension of latent space, and $\beta$ refers to the smoothness index of the latent density function. This rate mitigates the curse of dimensionality inherent in raw data, thereby illustrating the effectiveness of latent diffusion models from a theoretical perspective. Moreover, the framework of our theoretical analysis is also applicable for analyzing other diffusion models.
@@ 1.1 Related work
Normalizing flows [DSDB16, RM15, PNR+21], especially those guided by ordinary differential equations (ODEs), provide a robust framework for transforming a simple probability distribution, such as a standard normal distribution, into a more complex form through a sequence of invertible and differentiable mappings. The essence of normalizing flows lies in their ability to capture complex dependencies and structures within data by iteratively applying invertible transformations. These transformations facilitate the modeling of sophisticated probability distributions, enabling practitioners to navigate through the intricacies of high-dimensional data and capture latent features. Recently, [CDD23, CCL+23] undertook a theoretical analysis of ODE flows, specifically introducing the accurate score estimation. However, this assumption warrants further verification. Very recently, [CDJ+24, GHJZ24, JLWY24] provided end-to-end convergence rates for ODE-based (conditional) generative models.
@@
In diffusion models, the fundamental architecture comprises both forward and backward processes. The primary objective of the forward process is to systematically add Gaussian noises to the dataset, thereby resulting in the limiting distribution being a Gaussian distribution. Conversely, the backward process assumes the role of a denoising mechanism, undertaking the formidable task of reconstructing data in a manner that closely approximates the characteristics of the original samples. Within the denoising step, a critical step emerges in the form of training a score neural network. This neural network is meticulously crafted to approximate the underlying score function, the gradient of the log density function of the underlying distribution. This synthesis of forward and backward processes, coupled with the sophisticated training of a score neural network, encapsulates the essence of diffusion models. Furthermore, within practical applications, datasets frequently exhibit inherent low-dimensional structures. The derivation of their low-dimensional representations within latent spaces is consequently pursued. A noteworthy advancement within this domain is the stable diffusion model, introduced by [RBL+22]. The integration of an autoencoder architecture within the stable diffusion model emerges as a pivotal mechanism, facilitating the discernment of the intrinsic low-dimensional essence of raw data. By executing diffusion processes within the latent spaces, the model excels in encapsulating the underlying distribution. This capability is particularly advantageous within complex data distributions. In addition, analogous studies have investigated the effective capture of data distributions through the utilization of latent spaces. Reference can be made to [VKK21, RDN+22]. In summary, these diffusion models can be generally classified into three categories: denoising diffusion probabilistic models, score-based generative models, and score SDEs. A thorough examination of diffusion models can be found in the specialized review paper [YZS+23]. The theoretical analysis of diffusion models primarily focuses on error analysis with accurate score estimators [CLL23, CDS23, LLT22, LLT23, BDBDD23, LWCC23, GNZ23] or end-to-end error analysis [OAS23, CHZW23]. Recently, an end-to-end theoretical analysis of conditional diffusion models was explored in [FYWC24]. However, these theoretical results did not consider the theoretical analysis of pre-training encoder-decoder structures, a crucial step in contemporary mainstream diffusion models. Consequently, these findings are unable to provide further insights into the underlying mechanisms of diffusion models. In contrast, our theoretical results include the theoretical analysis of pre-training encoder-decoder architectures, thereby offering a comprehensive explanation.
@@ 1.2 Notations
We introduce the notations used throughout this paper. Let $[N]:=\{0,1,\cdots,N\}$ represent the set of integers ranging from 0 to $N$. Let $\mathbb{N}^{+}$ denote the set of positive integers. For matrices $A,B\in\mathbb{R}^{d\times d}$, we assert $A\preccurlyeq B$ when the matrix $B-A$ is positive semi-definite. The identity matrix in $\mathbb{R}^{d\times d}$ is denoted as $\mathbf{I}_{d}$. The $\ell^{2}$-norm of a vector $\mathbf{x}=\{x_{1},\ldots,x_{d}\}^{\top}\in\mathbb{R}^{d}$ is defined by $\|\mathbf{x}\|:=\sqrt{\sum_{i=1}^{d}x_{i}^{2}}$. Simultaneously, the operator norm of a matrix $A$ is articulated as $\|{A}\|:=\sup_{\|\mathbf{x}\|\leq{1}}\|{A\mathbf{x}}\|$. The function space $C^{2}(\mathbb{R}^{d})$ encompasses functions that are twice continuously differentiable from $\mathbb{R}^{d}$ to $\mathbb{R}$. For any $f\in C^{2}(\mathbb{R}^{d})$, the symbols $\nabla{f}$, $\nabla^{2}{f}$, and ${\rm{\Delta}}f$ signify its gradient, Hessian matrix, and Laplacian, respectively. The $L^{\infty}(K)$-norm, denoted as $\|{f}\|_{L^{\infty}(K)}:=\sup_{\mathbf{x}\in K}|f(\mathbf{x})|$, captures the supremum of the absolute values of a function over a set $K\subset\mathbb{R}^{d}$. For a vector function $\mathbf{v}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, the $L^{\infty}(K)$-norm is defined as $\|\mathbf{v}\|_{L^{\infty}(K)}:=\sup_{\mathbf{x}\in K}\|\mathbf{v}(\mathbf{x})\|$. The asymptotic notation $f(\mathbf{x})=\mathcal{O}\left(g(\mathbf{x})\right)$ is employed to signify that $f(\mathbf{x})\leq Cg(\mathbf{x})$ for some constant $C>0$. Additionally, the notation $\widetilde{\mathcal{O}}(\cdot)$ is utilized to discount logarithmic factors in the asymptotic analysis.
@@ 2 Preliminaries
A class of neural networks NN$(L,M,J,\kappa)$ with depth $L$, width $M$, sparsity level $J$, weight bound $\kappa$, is defined as
@@
$\mathbf{W}_{i}\in\mathbb{R}^{d_{i+1}\times d_{i}},\mathbf{b}_{i}\in\mathbb{R}^{d_{i+1}},i=0,1,\ldots,L-1,M:=\max\{d_{0},\ldots,d_{L}\},$
$\mathop{\max}_{1\leq{i}\leq{L}}\{\|{\mathbf{b}}_{i}\|_{\infty},\|{\mathbf{W}_{i}}\|_{\infty}\}\leq{\kappa},\leavevmode\nobreak\ \sum_{i=1}^{L}\left(\|{\mathbf{W}_{i}}\|_{0}+\|{\mathbf{b}_{i}}\|_{0}\right)\leq J\Big{\}}.$
@@
Definition 2.4 (($\beta$, $R$)-Hölder Class).
Let $\beta=r+\gamma>0$ be a degree of smoothness, where $r=\lfloor\beta\rfloor$ is an integer and $\gamma\in[0,1)$. For a function $f:\Omega\rightarrow\mathbb{R}$, its Hölder norm is defined as
$\|f\|_{\mathcal{H}^{\beta}}:=\max_{\|\boldsymbol{\alpha}\|_{1}\leq r}\|\partial^{\boldsymbol{\alpha}}f\|_{\infty}+\max_{\|\boldsymbol{\alpha}\|_{1}=r}\sup_{\mathbf{x}\neq\mathbf{y}}\frac{|\partial^{\boldsymbol{\alpha}}f(\mathbf{x})-\partial^{\boldsymbol{\alpha}}f(\mathbf{y})|}{\|\mathbf{x}-\mathbf{y}\|^{\gamma}},$
where $\boldsymbol{\alpha}$ is a multi-index. We say a function $f$ is $\beta$-Hölder, if and only if $\|f\|_{\mathcal{H}^{\beta}}<\infty$. The ($\beta$, $R$)-Hölder class for some constant $R>0$ is defined as
$\mathcal{H}^{\beta}(\Omega,R):=\Big{\{}f:\Omega\rightarrow\mathbb{R}\Big{|}\|f\|_{\mathcal{H}^{\beta}}\leq R\Big{\}}.$
@@ 3.1 Pre-training
We observe that the derivation of $(\widehat{\boldsymbol{E}},\widehat{\boldsymbol{D}})$ as defined in (3) can be facilitated through the utilization of pre-existing large-scale models. This approach permits accommodation for distributional shifts that may deviate from the target distribution. As a result, it guarantees an abundant supply of samples essential for pre-training endeavors, enabling us to obtain convergence rates that circumvent the curse of dimensionality, as discussed in our theoretical development in Subsection 4.2.3 and Section 4.3.
@@ 3.2 Latent Diffusion Modeling
With the trained encoder-decoder structure established in Section 3.1, we can subsequently construct the diffusion model in the latent space. We denote by $\widehat{\mathbf{x}}:=\widehat{\boldsymbol{E}}\mathbf{x}$ with $\mathbf{x}\sim p_{data}(\cdot)$. Let ${p}^{*}_{data}(\cdot)$ and $\widehat{p}^{*}_{data}(\cdot)$ represent the density functions of $\boldsymbol{E}^{*}\mathbf{x}$ and $\widehat{\mathbf{x}}$, respectively. Additionally, let $\Phi_{\sigma}(\cdot)$ denote the density of the normal distribution $\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{d^{*}})$. Let $\mathcal{X}:=\{\mathbf{x}_{i}\}_{i=1}^{n}$ be i.i.d. from $p_{data}$, then we have access to the encoder data $\widehat{\mathcal{X}}:=\{\widehat{\mathbf{x}}_{i}\}_{i=1}^{n}=\{\widehat{\boldsymbol{E}}\mathbf{x}_{i}\}_{i=1}^{n}$. We denote the convolution of $\widehat{p}^{*}_{data}$ and $\Phi_{\sigma}(\cdot)$ as $q(\cdot,\cdot)$, defined as follows:
@@
$\mathrm{d}\mathbf{x}_{t}=\sigma^{2}\nabla\log{q}_{1-t}(\mathbf{x}_{t})\mathrm{d}t+\sigma\mathrm{d}\mathbf{w}_{t},\leavevmode\nobreak\ t\in[0,1],\leavevmode\nobreak\ \mathbf{x}_{0}\sim{q}({\sigma},\mathbf{x}),$ (5)
where $q_{1-t}(\cdot):=q(\sqrt{1-t}\sigma,\cdot)$, $\mathbf{w}_{t}$ represents a $d^{*}$-dimensional Brownian motion. Consequently, $\mathbf{x}_{1}$ follows the distribution $\widehat{p}^{*}_{data}$ as established in Proposition 2.2 and [WJX+21, Theorem 3]. This formalization provides a coherent and rigorous foundation for our generative framework.
Score Matching. The SDE provided in (5) offers a pathway for constructing a generative model through the application of EM discretization, contingent upon access to the target score denoted as $\mathbf{s}^{*}(t,\mathbf{x}):=\nabla\log{q}_{t}(\mathbf{x})$. However, in practical scenarios, this score function remains elusive due to its intricate connection with the target density $\widehat{p}^{*}(\cdot)$. Consequently, we resort to score matching techniques [HD05, Vin11] employing deep ReLU FNNs.
Verification reveals that $\mathbf{s}^{*}$ minimizes the loss function $\mathcal{L}(\mathbf{s})$ across all measurable functions, with $\mathcal{L}(\mathbf{s})$ expressed as
$\mathcal{L}(\mathbf{s}):=\frac{1}{T}\int_{0}^{T}\left(\mathbb{E}_{\mathbf{x}_{t}\sim q_{1-t}(\mathbf{x})}\|{\mathbf{s}(1-t,\mathbf{x}_{t})-\nabla\log{q_{1-t}(\mathbf{x}_{t})}}\|^{2}\right)\mathrm{d}t,$
where $10<T<$ 1. We notice that the deliberate selection of $T$ is motivated by the necessity to preclude the score function from exhibiting a blow-up at $T=1$, concomitantly facilitating the stabilization of the training process for the model. A meticulous exposition regarding the determination of $T$ is provided in Subsection 4.2.3. Furthermore, the application of this time truncation technique extends beyond our work and has found utility in both model training and theoretical analyses in diffusion models [SE20, VKK21, CHZW23, CLL23, OAS23].